OpenStudy (anonymous):

can you tell me what formula in this questions x-2y=17 , 3x-4y=1

OpenStudy (anonymous):

Its simultaneous equations right? so multiply the first equation (1) x2 to get 2x-4y=34. Call it equation (3). Then do equation (2) - equation (3) which gives x = - 33 Then sub into equation (1) -33 -2y = 17 2y = 17+33 = 50 y = 25

OpenStudy (anonymous):

whooops y = -25 my bad

OpenStudy (anonymous):

nevermine , thank you